INAF queue observing, 2014 Feb 02-03

Observer: E. Sani, F. Cusano, S. Bisogni, R. Carini
Telescope Operator: D. Gionzales Huerta
Support Astronomer: O. Kuhn on call
Instrument: MODS

Summary

At the beginning of the night we stay close because of dewpoint reaching ext temperature.
While we are waiting the TO starts the mirror ventilation.
At the beginning of scientific operations there are clouds passing, 60% humidity and no wind.
Since there are thick clouds we can't observe scientific targets: no collimation (the WFS can't sent corrections to collimate) + no guiding (the GCS can't see a R=12 star).
We try to collimate on a very bright star, this does not work and the TO closes the dome.
We stay closed the rest of the night.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Standard Feige 34 red grating - 1.1

Details

03:40 The TO opens the dome
  • Still too cloudy we are no able to collimate. We close again.

* *Dome opened

08:09 Slew to the standard Feige 34
  • execMODS feige34.acq
  • modsAlign -r mods1r 0004

08:14 execMODS feige34.obs mods1r 0005-0007

08:28 acqMODS 4Q1_J1048_M1_mos.acq mods1r 0008-00010
  • modsAlign mods.6.mms mods1r 008-0009

08:45 execMODS 4Q1_J1048_M1_mos_5.obs mods1r 0011-00012
  • at the beginning of the integration the seeing is 1.8-2.0 on the GCS
  • guide star lost due to thick clouds. We have to wait the weather clears.

09:08 *The TO closes the dome *

-- EleanoraSani - 03 Feb 2014
Topic revision: r5 - 04 Feb 2014, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback